# The Cofinality of Generating Families

### v1

@@ 4.2 How Many Compact Sets Suffice?
Our results above on $\mathop{seq}(M)$ - building on the characterization of $\mathbf{seq}(M)$ up to Tukey equivalence obtained in [7], along with our results on the sampling number - tell us the following: that $\mathop{seq}(M)$ is ranked by the size of $M$, if $|M|\leq|N|$ then $\mathop{seq}(M)\leq\mathop{seq}(N)$, and for a fixed size of space at most two values of $\mathop{seq}(M)$ can occur: for every uncountable $\kappa\leq\mathfrak{c}$
$\{\mathop{seq}(M):|M|=\kappa\}\subseteq\left\{\lim_{\mu<|M|}\mathop{sam}(\mu)\cdot\mathfrak{b},\mathop{sam}(|M|)\cdot\mathfrak{b}\right\}.$
@@
Nevertheless we show below that $kc(M)$ plays a role for $k(M)$ loosely analogous to the role that $|M|$ plays for $\mathop{seq}(M)$, with the connection being made by the sampling number combined with the bounding number. Theorem 4.7 below implies that: for every uncountable $\kappa\leq\mathfrak{c}$
$\{\mathop{k}(M):kc(M)=\kappa\}\subseteq[\kappa\cdot\mathfrak{b},\mathop{sam}(\kappa)\cdot\mathfrak{b}].$
Above we explored, in the context of $\mathop{seq}(M)$, the possible values of $\mathop{sam}(\kappa)$ and the gap between $\kappa$ and $\mathop{sam}(\kappa)$, for $\kappa=|M|$. We do not rehash here the situation when $\kappa=kc(M)$. Instead in the sequel we focus on what is different between $\mathop{seq}(M)$ and $k(M)$, namely (1) that while $\mathop{seq}(M)$ for $|M|=\kappa$ can only take on the value of one of a lower and an upper bound, $\{\mathop{k}(M):kc(M)=\kappa\}$ is constrained merely to an interval, and (2) the lower bound ($\kappa\cdot\mathfrak{b}$) on this interval of $k(M)$ values is potentially smaller than the analogous lower bound value for $\mathop{seq}(M)$ values ($\lim_{\mu<\kappa}\mathop{sam}(\mu)\cdot\mathfrak{b}$). Our task is made more difficult because we have no analog to the realization of types theorem, Theorem 4.2, for $\mathbf{seq}(M)$.
Let $\kappa$ be an uncountable cardinal with $\kappa\leq\mathfrak{c}$. Define the $k$-spread at $\kappa$, to be $\sigma_{k}(\kappa)=|\{k(M):kc(M)=\kappa\}|$. We show in Theorem 4.8 that the upper bound of the interval, $\kappa\cdot\mathfrak{b},\mathop{sam}(\kappa)\cdot\mathfrak{b}]$, is always attained; that is, for some $M$ we have $kc(M)=\kappa$ and $k(M)=\mathop{sam}(\kappa)\cdot\mathfrak{b}$, and hence $\sigma_{k}(\kappa)\geq 1$. We then show, Theorem 4.11, that in certain circumstances the lower bound, $\kappa\cdot\mathfrak{b}$, is also attained. It follows that $\sigma_{k}(\kappa)$ may be at least two. Finally we show below that, modulo large cardinals, for every $2\leq K<\omega$ it is consistent that for some $\kappa$ we have $\sigma_{k}(\kappa)=K$ (Theorem 4.12) and it is consistent that for some $\kappa$ we have $\sigma_{k}(\kappa)\geq\aleph_{0}$ (Theorem 4.13).
@@
We now see that for non $\sigma$-compact separable metrizable spaces the $k$-ness number is connected to the compact-covering number via the sampling number.
@@
$\mathop{sam}(kc(M))\cdot\mathfrak{b}\geq k(M)\geq kc(M)\cdot\mathfrak{b}.$
@@
A separable metrizable space $M$ is totally imperfect if every compact subset of $M$ is countable, equivalently - since uncountable compact metrizable spaces contain a Cantor set - if $M$ contains no compact subsets of size continuum. Note that if $M$ has size strictly less than the continuum then $M$ is totally imperfect; in particular, locally small spaces are totally imperfect (if $M$ is locally small then $\operatorname{cof}(|M|)=\omega$, but $\operatorname{cof}(\mathfrak{c})\neq\omega$). However there are totally imperfect separable metrizable spaces of size continuum, Bernstein subsets of the reals, for example. We can determine the $k$-ness number of totally imperfect spaces exactly in terms of $\mathop{sam}(kc(M))$ and the bounding number.
@@
    If $M$ is totally imperfect but not locally small then $k(M)=\mathop{sam}(kc(M))\cdot\mathfrak{b}$.
@@
    If $M$ is locally small then $k(M)=\lim_{\mu<kc(M)}\mathop{sam}(\mu)\cdot\mathfrak{b}$.
@@
It follows from (i) that the top of the range, $\mathop{sam}(\kappa)\cdot\mathfrak{b}$, is always obtained, in ZFC.
@@
If $\omega<\kappa\leq\mathfrak{c}$, then there is a separable metrizable space $M_{\kappa}$ so that $k(M_{\kappa})=\mathop{sam}(\kappa)\cdot\mathfrak{b}$.
@@
If $\omega<\kappa\leq\mathfrak{c}$ and $\operatorname{cof}(\kappa)=\omega$, then there is a separable metrizable space $M_{\kappa}$ so that $k(M_{\kappa})=\lim_{\mu<\kappa}\mathop{sam}(\mu)\cdot\mathfrak{b}$.
This may, or may not, place an additional value in $\{k(M):\mathop{kc}(M)=\kappa\}$ - depending on whether $\lim_{\mu<\kappa}\mathop{sam}(\mu)\cdot\mathfrak{b}$ and $\mathop{sam}(\kappa)\cdot\mathfrak{b}$ are equal.
@@
A natural question is whether, in the above scenario where $M=N^{\omega}$, they have the same $k$-ness numbers: if $N$ is non-compact, separable metrizable then is $k(N^{\omega})=k(N)$? This is connected to Question C. Indeed, if there is a cardinal $\kappa$ with $\mathfrak{d}\leq\kappa<\mathfrak{c}$ and $\mathop{sam}(\kappa)<\mathop{cov}(\kappa)$ then the answer to the above question is in the negative. To see this, let $N$ be a subset of the reals with $|N|=\kappa$. Then $k(N^{\omega})=\mathop{cov}(|N|)>\mathop{sam}(|N|)=k(N)$.
@@
Recall that the $k$-spread at an uncountable $\kappa\leq\mathfrak{c}$ is $\sigma_{k}(\kappa)=|\{k(M):kc(M)=\kappa\}|$, namely, the number of values of $k$ attained for a fixed $kc$. We know from Theorem 4.8(i) that we always have $\sigma_{k}(\kappa)\geq 1$. Further, for every $n\in\mathbb{N}$ we know $\mathop{sam}(\aleph_{n})=\aleph_{n}$, so $\sigma_{k}(\aleph_{n})=1$. We now show that, modulo large cardinals, for any finite $K\geq 2$ there exist a model and a $\kappa$ with $\sigma_{k}(\kappa)=K$, and also a model and a $\kappa$ with $\sigma_{k}(\kappa)\geq\aleph_{0}$.
@@
Modulo large cardinals, it is consistent that $\mathfrak{b}=\aleph_{1}$, $\mathop{sam}(\lambda^{+})=\lambda^{+K}$, and for $n=1,\ldots,K$ there are separable metrizable spaces, $M_{n}$, such that $kc(M_{n})=\lambda^{+}$ while $k(M_{n})=\lambda^{+n}$.
@@
$\{k(M):kc(M)=\kappa\}=[\lambda^{+},\lambda^{+K}]=[\kappa\cdot\mathfrak{b},\mathop{sam}(\kappa)\cdot\mathfrak{b}],\ \text{ so }\sigma_{k}(\kappa)=K.$
@@
Modulo large cardinals, it is consistent that $\mathfrak{b}=\aleph_{1}$, $\mathop{sam}(\lambda^{+})\geq\lambda^{+\omega}$, and for each integer $n\geq 1$ there are separable metrizable spaces, $M_{n}$, such that $kc(M_{n})=\lambda^{+}$ while $k(M_{n})=\lambda^{+n}$.
@@
$\lambda^{+},\lambda^{+\omega})\subseteq\{k(M):kc(M)=\kappa\}\subseteq[\kappa\cdot\mathfrak{b},\mathop{sam}(\kappa)\cdot\mathfrak{b}],\ \text{ so }\sigma_{k}(\kappa)\geq\aleph_{0}.$
In order to prove both these results we need to first construct, with the aid of large cardinals, models with suitable configurations of sampling numbers, and then to find within those models the desired spaces, $M_{n}$. For the first step we rely on recent work of the second author and Cummings, [3].
@@
Fix $K\geq 2$. Start with the model of Gilton and Cummings above. For each $1\leq n\leq K$, apply (3) together with Corollary 3.9 to conclude that $\mathop{sam}(\aleph_{\omega\cdot n})\geq\aleph_{\omega\cdot K+n}$. By (2), $\mathop{cov}(\aleph_{\omega\cdot n})\leq\aleph_{\omega\cdot K+n}$, whence $\mathop{sam}(\aleph_{\omega\cdot n})$ and $\mathop{cov}(\aleph_{\omega\cdot n})$ are both equal to $\aleph_{\omega\cdot K+n}$.
Now do random real forcing to increase the value of $\mathfrak{c}$ strictly above $\aleph_{\omega\cdot K+K}$, while keeping $\mathfrak{b}$ and $\mathfrak{d}$ fixed at $\omega_{1}$. By Lemmas 3.14 and 3.13, this forcing doesn’t change the relevant values of the sampling and covering numbers. Thus, in the resulting model, (a) $\mathfrak{b}=\mathfrak{d}=\omega_{1}$; (b) $\mathfrak{c}>\aleph_{\omega\cdot K+K}$; and (c) for each $1\leq n\leq K$, $\mathop{cov}(\aleph_{\omega\cdot n})=\mathop{sam}(\aleph_{\omega\cdot n})=\aleph_{\omega\cdot K+n}$.
@@
$\mathop{sam}(\lambda^{+})=\mathop{sam}(\aleph_{\omega\cdot K+1})=\mathop{sam}(\aleph_{\omega\cdot K})\cdot\aleph_{\omega\cdot K+1}=\aleph_{\omega\cdot K+K}=\lambda^{+K}.$
@@
The $k$-ness number of $M_{n}$, $k(M_{n})$, is the larger of $k(M_{\infty})$ and $k(N_{n})$. From Theorem 4.11 we know $k(M_{\infty})=kc(M_{\infty})=\aleph_{\omega\cdot K+1}$. While from Theorem 4.8, as $kc(N_{n})=\aleph_{\omega\cdot n}<\mathfrak{c}$ (by (b)), we have $k(N_{n})=\mathop{sam}(\aleph_{\omega\cdot n})=\aleph_{\omega\cdot K+n}$. Hence $k(M_{n})=\aleph_{\omega\cdot K+n}=\lambda^{+n}$, as required. ∎
@@
The argument is similar to that for Theorem 4.12 and we omit some details. Start with the model of Gilton and Cummings above. Applying a suitable random real forcing we get a model where (a) $\mathfrak{b}=\omega_{1}$, (b) $\mathfrak{c}>\aleph_{\omega^{2}+\omega}$, and (c) $\mathop{cov}(\aleph_{\omega\cdot n})=\mathop{sam}(\aleph_{\omega\cdot n})=\aleph_{\omega^{2}+n}$, for every $n\geq 1$.
@@
There is much that we do not know about the $k$-spread and the set $\{k(M):\mathop{kc}(M)=\kappa\}$. We know, from the preceding result, it is consistent that for $\kappa=\aleph_{\omega^{2}+1}$ the $k$-spread can be at least $\aleph_{0}$. Can the $k$-spread of some uncountable $\kappa\leq\mathfrak{c}$ be uncountable? Can $\{k(M):\mathop{kc}(M)=\kappa\}$ be a proper subset of $\kappa\cdot\mathfrak{b},\mathop{sam}(\kappa)\cdot\mathfrak{b}]$?
@@
For the second claim we work in a model where $\aleph_{1}<\mathfrak{p}\leq\mathfrak{b}<\mathfrak{d}$, and note that this is consistent. Take any $\aleph_{1}$-sized subset, $M$ say, of the reals. Then $M$ is co-analytic because $|M|=\aleph_{1}<\mathfrak{p}$. But $M$ is not $\sigma$-compact, in fact $\mathop{kc}(M)=\aleph_{1}$, because $|M|=\aleph_{1}<\mathfrak{c}$. Hence $\mathop{sam}(kc(M))\cdot\mathfrak{b}=\aleph_{1}\cdot\mathfrak{b}=\mathfrak{b}$ (Lemma 3.3), so $\mathop{k}(M)=\mathfrak{b}$ (Theorem 4.7) and $\mathfrak{b}<\mathfrak{d}$ in our model. ∎

### v2

@@ 3.2 How Many Compact Sets Suffice?
Our results above on $\mathop{seq}(M)$ - building on the characterization of $\mathbf{seq}(M)$ up to Tukey equivalence obtained in [8], along with what is known about the covering number - tell us the following: that $\mathop{seq}(M)$ is ranked by the size of $M$, if $|M|\leq|N|$ then $\mathop{seq}(M)\leq\mathop{seq}(N)$, and for a fixed size of space at most two values of $\mathop{seq}(M)$ can occur: for every uncountable $\kappa\leq\mathfrak{c}$
$\{\mathop{seq}(M):|M|=\kappa\}\subseteq\left\{\lim_{\mu<|M|}\mathop{cov}(\mu)\cdot\mathfrak{b},\mathop{cov}(|M|)\cdot\mathfrak{b}\right\}.$
@@
Nevertheless we show below that $kc(M)$ plays a role for $k(M)$ loosely analogous to the role that $|M|$ plays for $\mathop{seq}(M)$, with the connection being made by the covering number combined with the bounding number. Theorem 3.7 below implies that: for every uncountable $\kappa\leq\mathfrak{c}$
$\{\mathop{k}(M):kc(M)=\kappa\}\subseteq[\kappa\cdot\mathfrak{b},\mathop{cov}(\kappa)\cdot\mathfrak{b}].$
Above we explored, in the context of $\mathop{seq}(M)$, the possible values of $\mathop{cov}(\kappa)$ and the gap between $\kappa$ and $\mathop{cov}(\kappa)$, for $\kappa=|M|$. We do not rehash here the situation when $\kappa=kc(M)$. Instead in the sequel we focus on what is different between $\mathop{seq}(M)$ and $k(M)$, namely (1) that while $\mathop{seq}(M)$ for $|M|=\kappa$ can only take on the value of one of a lower and an upper bound, $\{\mathop{k}(M):kc(M)=\kappa\}$ is constrained merely to an interval, and (2) the lower bound ($\kappa\cdot\mathfrak{b}$) on this interval of $k(M)$ values is potentially smaller than the analogous lower bound value for $\mathop{seq}(M)$ values ($\lim_{\mu<\kappa}\mathop{cov}(\mu)\cdot\mathfrak{b}$). Our task is made more difficult because we have no analog to the realization of types theorem, Theorem 3.2, for $\mathbf{seq}(M)$.
Let $\kappa$ be an uncountable cardinal with $\kappa\leq\mathfrak{c}$. Define the $k$-spread at $\kappa$, to be $\sigma_{k}(\kappa)=|\{k(M):kc(M)=\kappa\}|$. We show in Theorem 3.8 that the upper bound of the interval, $\kappa\cdot\mathfrak{b},\mathop{cov}(\kappa)\cdot\mathfrak{b}]$, is always attained; that is, for some $M$ we have $kc(M)=\kappa$ and $k(M)=\mathop{cov}(\kappa)\cdot\mathfrak{b}$, and hence $\sigma_{k}(\kappa)\geq 1$. We then show, Theorem 3.11, that in certain circumstances the lower bound, $\kappa\cdot\mathfrak{b}$, is also attained. It follows that $\sigma_{k}(\kappa)$ may be at least two. Finally we show below that, modulo large cardinals, for every $2\leq K<\omega$ it is consistent that for some $\kappa$ we have $\sigma_{k}(\kappa)=K$ (Theorem 3.12) and it is consistent that for some $\kappa$ we have $\sigma_{k}(\kappa)\geq\aleph_{0}$ (Theorem 3.13).
@@
We now see that for non $\sigma$-compact separable metrizable spaces the $k$-ness number is connected to the compact-covering number via the covering number.
@@
$\mathop{cov}(kc(M))\cdot\mathfrak{b}\geq k(M)\geq kc(M)\cdot\mathfrak{b}.$
@@
A separable metrizable space $M$ is totally imperfect if every compact subset of $M$ is countable, equivalently - since uncountable compact metrizable spaces contain a Cantor set - if $M$ contains no compact subsets of size continuum. Note that if $M$ has size strictly less than the continuum then $M$ is totally imperfect; in particular, locally small spaces are totally imperfect (if $M$ is locally small then $\operatorname{cof}(|M|)=\omega$, but $\operatorname{cof}(\mathfrak{c})\neq\omega$). However there are totally imperfect separable metrizable spaces of size continuum, Bernstein subsets of the reals, for example. We can determine the $k$-ness number of totally imperfect spaces exactly in terms of $\mathop{cov}(kc(M))$ and the bounding number.
@@
    If $M$ is totally imperfect but not locally small then $k(M)=\mathop{cov}(kc(M))\cdot\mathfrak{b}$.
@@
    If $M$ is locally small then $k(M)=\lim_{\mu<kc(M)}\mathop{cov}(\mu)\cdot\mathfrak{b}$.
@@
It follows from (i) that the top of the range, $\mathop{cov}(\kappa)\cdot\mathfrak{b}$, is always obtained, in ZFC.
@@
If $\omega<\kappa\leq\mathfrak{c}$, then there is a separable metrizable space $M_{\kappa}$ so that $k(M_{\kappa})=\mathop{cov}(\kappa)\cdot\mathfrak{b}$.
@@
If $\omega<\kappa\leq\mathfrak{c}$ and $\operatorname{cof}(\kappa)=\omega$, then there is a separable metrizable space $M_{\kappa}$ so that $k(M_{\kappa})=\lim_{\mu<\kappa}\mathop{cov}(\mu)\cdot\mathfrak{b}$.
This may, or may not, place an additional value in $\{k(M):\mathop{kc}(M)=\kappa\}$ - depending on whether $\lim_{\mu<\kappa}\mathop{cov}(\mu)\cdot\mathfrak{b}$ and $\mathop{cov}(\kappa)\cdot\mathfrak{b}$ are equal.
@@
A natural question is whether, in the above scenario where $M=N^{\omega}$, they have the same $k$-ness numbers: if $N$ is non-compact, separable metrizable then is $k(N^{\omega})=k(N)$?
@@
Recall that the $k$-spread at an uncountable $\kappa\leq\mathfrak{c}$ is $\sigma_{k}(\kappa)=|\{k(M):kc(M)=\kappa\}|$, namely, the number of values of $k$ attained for a fixed $kc$. We know from Theorem 3.8(i) that we always have $\sigma_{k}(\kappa)\geq 1$. Further, for every $n\in\mathbb{N}$ we know $\mathop{cov}(\aleph_{n})=\aleph_{n}$, so $\sigma_{k}(\aleph_{n})=1$. We now show that, modulo large cardinals, for any finite $K\geq 2$ there exist a model and a $\kappa$ with $\sigma_{k}(\kappa)=K$, and also a model and a $\kappa$ with $\sigma_{k}(\kappa)\geq\aleph_{0}$.
@@
Modulo large cardinals, it is consistent that $\mathfrak{b}=\aleph_{1}$, $\mathop{cov}(\lambda^{+})=\lambda^{+K}$, and for $n=1,\ldots,K$ there are separable metrizable spaces, $M_{n}$, such that $kc(M_{n})=\lambda^{+}$ while $k(M_{n})=\lambda^{+n}$.
@@
$\{k(M):kc(M)=\kappa\}=[\lambda^{+},\lambda^{+K}]=[\kappa\cdot\mathfrak{b},\mathop{cov}(\kappa)\cdot\mathfrak{b}],\ \text{ so }\sigma_{k}(\kappa)=K.$
@@
Modulo large cardinals, it is consistent that $\mathfrak{b}=\aleph_{1}$, $\mathop{cov}(\lambda^{+})\geq\lambda^{+\omega}$, and for each integer $n\geq 1$ there are separable metrizable spaces, $M_{n}$, such that $kc(M_{n})=\lambda^{+}$ while $k(M_{n})=\lambda^{+n}$.
@@
$\lambda^{+},\lambda^{+\omega})\subseteq\{k(M):kc(M)=\kappa\}\subseteq[\kappa\cdot\mathfrak{b},\mathop{cov}(\kappa)\cdot\mathfrak{b}],\ \text{ so }\sigma_{k}(\kappa)\geq\aleph_{0}.$
In order to prove both these results we need to first construct, with the aid of large cardinals, models with suitable configurations of covering numbers, and then to find within those models the desired spaces, $M_{n}$. For the first step we rely on recent work of the second author and Cummings, [4].
@@
Fix $K\geq 2$. Start with the model of Gilton and Cummings above. For each $1\leq n\leq K$, from (2), (3) together with Corollary 2.15 conclude that $\mathop{cov}(\aleph_{\omega\cdot n})$ is equal to $\aleph_{\omega\cdot K+n}$. Now do random real forcing to increase the value of $\mathfrak{c}$ strictly above $\aleph_{\omega\cdot K+K}$, while keeping $\mathfrak{b}$ and $\mathfrak{d}$ fixed at $\omega_{1}$. By Lemma 2.9, this forcing doesn’t change the relevant values of the covering numbers. Thus, in the resulting model, (a) $\mathfrak{b}=\mathfrak{d}=\omega_{1}$; (b) $\mathfrak{c}>\aleph_{\omega\cdot K+K}$; and (c) for each $1\leq n\leq K$, $\mathop{cov}(\aleph_{\omega\cdot n})=\aleph_{\omega\cdot K+n}$.
@@
$\mathop{cov}(\lambda^{+})=\mathop{cov}(\aleph_{\omega\cdot K+1})=\mathop{cov}(\aleph_{\omega\cdot K})\cdot\aleph_{\omega\cdot K+1}=\aleph_{\omega\cdot K+K}=\lambda^{+K}.$
@@
The $k$-ness number of $M_{n}$, $k(M_{n})$, is the larger of $k(M_{\infty})$ and $k(N_{n})$. From Theorem 3.11 we know $k(M_{\infty})=kc(M_{\infty})=\aleph_{\omega\cdot K+1}$. While from Theorem 3.8, as $kc(N_{n})=\aleph_{\omega\cdot n}<\mathfrak{c}$ (by (b)), we have $k(N_{n})=\mathop{cov}(\aleph_{\omega\cdot n})=\aleph_{\omega\cdot K+n}$. Hence $k(M_{n})=\aleph_{\omega\cdot K+n}=\lambda^{+n}$, as required. ∎
@@
The argument is similar to that for Theorem 3.12 and we omit some details. Start with the model of Gilton and Cummings above. Applying a suitable random real forcing we get a model where (a) $\mathfrak{b}=\omega_{1}$, (b) $\mathfrak{c}>\aleph_{\omega^{2}+\omega}$, and (c) $\mathop{cov}(\aleph_{\omega\cdot n})=\aleph_{\omega^{2}+n}$, for every $n\geq 1$.
@@
There is much that we do not know about the $k$-spread and the set $\{k(M):\mathop{kc}(M)=\kappa\}$. We know, from the preceding result, it is consistent that for $\kappa=\aleph_{\omega^{2}+1}$ the $k$-spread can be at least $\aleph_{0}$. Can the $k$-spread of some uncountable $\kappa\leq\mathfrak{c}$ be uncountable? Can $\{k(M):\mathop{kc}(M)=\kappa\}$ be a proper subset of $\kappa\cdot\mathfrak{b},\mathop{cov}(\kappa)\cdot\mathfrak{b}]$?
@@
For the second claim we work in a model where $\aleph_{1}<\mathfrak{p}\leq\mathfrak{b}<\mathfrak{d}$, and note that this is consistent. Take any $\aleph_{1}$-sized subset, $M$ say, of the reals. Then $M$ is co-analytic because $|M|=\aleph_{1}<\mathfrak{p}$. But $M$ is not $\sigma$-compact, in fact $\mathop{kc}(M)=\aleph_{1}$, because $|M|=\aleph_{1}<\mathfrak{c}$. Hence $\mathop{cov}(kc(M))\cdot\mathfrak{b}=\aleph_{1}\cdot\mathfrak{b}=\mathfrak{b}$ (Lemma 2.4), so $\mathop{k}(M)=\mathfrak{b}$ (Theorem 3.7) and $\mathfrak{b}<\mathfrak{d}$ in our model. ∎
4 Acknowledgements
Thanks to Will Brian for his interest in this work, and for informing us [3] about the equivalence of the sampling and covering numbers.